OpenStudy (anonymous):

Fan and Medal! Elizabeth is hired to plan a class reunion. There are 275 students, but 256 of them will be attending. The school color is blue and white. There are 84 blue balloons and 96 white balloons. How may balloons bouquets can she make if she wants to have the same number each color in each bouquet and use all of her balloons?

OpenStudy (anonymous):

@bohotness

OpenStudy (bohotness):

yes

OpenStudy (anonymous):

help! @bohotness

OpenStudy (paxpolaris):

i am thinking this has something to do with GCF ...

OpenStudy (anonymous):

i dont even know @PaxPolaris

OpenStudy (paxpolaris):

84 = 12*7 96 = 12*8 GCF is 12

OpenStudy (paxpolaris):

so you make 12 bouquets ... each with 7 blue balloons and 8 black.

OpenStudy (anonymous):

it said that you ahve to have the equal number of color balloons in each bouquet... @PaxPolaris

OpenStudy (anonymous):

@PaxPolaris

OpenStudy (anonymous):

help

OpenStudy (paxpolaris):

i think the question is poorly worded ... i think what they are trying to say is that all bouquets have to be identical.

OpenStudy (anonymous):

oh ok. ill ask my teacher

OpenStudy (paxpolaris):

it is not possible for the bouquets to have the same number of blue and red balloons ... AND use up all the balloons

OpenStudy (anonymous):

oh. well, the book is stupid, it should go to hell and die in hellfire :D

OpenStudy (paxpolaris):

sorry .... lol i mean white when i said black or red

OpenStudy (anonymous):

yea i saw dat, just didnt feel like correcting a Mathlete

OpenStudy (anonymous):

XD